OpenStudy (anonymous):

At a farm stand, red peppers cost $4 per pound and green peppers cost $2 per pound. Tao wants to buy at least 8 pounds of peppers but spend no more than $28. The system of inequalities represents the number of pounds of red peppers, r, and the number of pounds of green peppers, g, Tao can buy. r + g ≥ 8 4r + 2g ≤ 28

OpenStudy (anonymous):

What are the minimum and maximum numbers of pounds of red peppers he can buy? The minimum is 0 pounds, and the maximum is 2 pounds. The minimum is 0 pounds, and the maximum is 6 pounds. The minimum is 2 pounds, and the maximum is 6 pounds. The minimum is 2 pounds, and the maximum is 7 pounds.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@iambatman

OpenStudy (mathstudent55):

\(r \ge -g + 8\) \(r \le -\dfrac{1}{2}g + 7\)

OpenStudy (mathstudent55):